OpenStudy (anonymous):

HELP WILL REWARD! What is the period of the function? y = 4 cos pi x A. 1 B. 2 C. pi D. 2pi

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

afk for a bit k

OpenStudy (anonymous):

back

OpenStudy (anonymous):

okay i'll be doing it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hmm its not coming to be clearly All i'am getting is this cos(pi.x) = 1 when x = 0, as cos(pi.0) = cos(0) = 1

OpenStudy (anonymous):

y=4sin(pi x) 4 2 2pi 2/pi 3

OpenStudy (anonymous):

Recall that the period for y = a*cos(bx + c) + d is given by 2π/|b|. Since y = cos(πx) = 1*cos(πx + 0) + 0, we see that b = π. Therefore, the period is: 2π/|b| = 2π/|π| = 2π/(π) = 2. I hope this helps, if not go to the link if you didn't recall anything ok.